# Abundant Population of Broad H Emitters in the GOODS-N Field Revealed by CONGRESS, FRESCO, and JADES

### v1

@@ Abstract
We present a spectroscopic search for broad H emitters at āā3.7ā6.5 in the GOODS-N field, utilizing JWST/NIRCam slitless spectroscopy from FRESCO and CONGRESS, complemented by JADES imaging. We identify 19 broad H emitters with FWHM ā1000 km/s at āā4ā5.5, including 9 new sources. The black hole masses and AGN bolometric luminosities, inferred from the broad H components, indicate that most sources are accreting at 10% of the Eddington limit. We derive their host stellar masses via SED fitting and find higher ratios relative to the local ā relations, consistent with previous studies. We find that 42% of the sample do not satisfy the widely-used color selection criteria for Little Red Dots (LRDs), with the majority of these sources lacking the characteristic steep red slope. A comparison of the average SEDs between our sample and LRDs selected in the same field reveals that the steep red slopes observed in some LRDs are likely due to line-boosting effects as previously suggested. Furthermore, we find that 68% of color-selected LRDs with H detections in the NIRCam/Grism spectra do not exhibit broad-line features. While the limited sensitivity of the grism spectra may hinder the detection of broad-line components in faint sources, our findings still highlight the enigmatic nature of the LRD population.
1 Introduction
@@
Supermassive black holes (SMBHs) are thought to reside at the centers of most massive galaxies (Kormendy & Ho, 2013; Greene etĀ al., 2020). The efficient accretion processes associated with these SMBHs contribute to the formation of active galactic nuclei (AGN), causing their host galaxies to exhibit broadened emission line profiles compared to non-active galaxies. Understanding the characteristics of AGN and accurately estimating their abundance in the early Universe () are essential for addressing key questions in extragalactic astronomy, including the reionization of the Universe and the co-evolution of AGN and their host galaxies.
@@
Prior to the launch of the JWST (Gardner etĀ al., 2023), studying the abundance and properties of faint AGN at high redshifts was highly challenging due to observational limitations. For AGN with ultraviolet (UV) magnitudes fainter than , determining their space density and physical properties was particularly difficult, with significant discrepancies reported (e.g., Parsa etĀ al., 2018; Giallongo etĀ al., 2019; Morishita etĀ al., 2020; Shen etĀ al., 2020; Finkelstein & Bagley, 2022). The advent of JWST, with its unparalleled infrared sensitivity, has opened new avenues for investigating these faint, distant objects. Recent JWST observations have demonstrated its remarkable capability to identify UV-faint AGN in the early Universe through various methodologies, including the detection of broad Balmer lines (Kocevski etĀ al., 2023; Harikane etĀ al., 2023; Oesch etĀ al., 2023; Larson etĀ al., 2023; Ćbler etĀ al., 2023; Barro etĀ al., 2024; Matthee etĀ al., 2024; Maiolino etĀ al., 2024; Greene etĀ al., 2024; Lin etĀ al., 2025a).
@@
Among the UV-faint AGN detected by JWST, an intriguing subset of the AGN population consists of sources with notably compact sizes and red colors. Matthee etĀ al. (2024) assembled one of the earliest samples of these sources, identified through broad H emission lines, and named them as ālittle red dotsā (LRDs). These LRDs appear as point sources and exhibit distinctive spectral energy distributions (SEDs), characterized by blue rest-frame UV slopes and extremely steep red rest-frame optical slopes (e.g., Greene etĀ al., 2024; LabbĆ© etĀ al., 2025). Such a āV-shapedā SED is found to be rare for AGN at 2 (e.g., Noboriguchi etĀ al. 2019). To investigate the origin of the UV and optical emission in LRDs, several recent studies have assembled samples of these objects by identifying compact sources and targeting those with āV-shapedā SEDs, using NIRCam color cuts as a selection criterion (Akins etĀ al., 2023; Barro etĀ al., 2024; Greene etĀ al., 2024; PĆ©rez-GonzĆ”lez etĀ al., 2024; Kokorev etĀ al., 2024; Williams etĀ al., 2024; Rinaldi etĀ al., 2024; LabbĆ© etĀ al., 2025).
@@
Although numerous photometrically selected LRD candidates are assembled, their true nature remains uncertain. It is still unclear whether these LRDs are primarily AGN host galaxies or if their āV-shapedā SEDs may instead be driven by star formation (Barro etĀ al., 2024; Akins etĀ al., 2023; Maiolino etĀ al., 2024). Greene etĀ al. (2024) obtained follow-up NIRSpec spectroscopy on 15 LRDs photometrically selected in LabbĆ© etĀ al. (2025) and found that 60% exhibited definitive evidence of broad H lines, confirming their AGN nature for a significant subset. In contrast, PĆ©rez-GonzĆ”lez etĀ al. (2024) reported that only 17% of their photometrically selected LRDs with available spectroscopy displayed broad line features. However, this fraction may be underestimated, as many existing broad lines could not be confirmed due to insufficient S/N of the grism spectra (Williams etĀ al., 2024). In Rinaldi etĀ al. (2024), analysis of the emission line properties of the LRD sample using NIRSpec spectra and classical line ratio diagnostics was unable to conclusively determine their nature, suggesting a likely mixed origin.
@@
In this work, we utilized grism spectroscopic data from the First Reionization Epoch Spectroscopically Complete Observations (FRESCO-PID: 1895; PI: P. Oesch; Oesch etĀ al. 2023) and the Complete NIRCam Grism Redshift Survey (CONGRESS-PID: 3577; PI: E. Egami; Sun et al., in prep, Lin etĀ al. 2025b), along with imaging data from the JWST Advanced Deep Extragalactic Survey (JADES-PID: 1181; PI: D. Eisenstein; Eisenstein etĀ al. 2023), to conduct a comprehensive search for broad H emitters at in the GOODS-North field (Giavalisco etĀ al. 2004, hereafter GOODS-N). After identifying this sample of broad H emitters, we aim to characterize their physical properties and compare them with LRDs recently identified in other studies. The environments of these AGNs and their impact on BH growth are discussed in a companion paper (Lin et al. in prep).
@@
This paper is organized as follows: In SectionĀ 2, we provide a brief overview of the spectroscopic and imaging data utilized in this study. SectionĀ 3 details the selection process for broad H emitters and the methodology for 1D line fitting. In SectionĀ 4, we present and discuss the results of the 1D line fitting for our broad H emitters. SectionĀ 5 explores the physical properties of the broad H emitter sample and compares our sample with the LRD sample identified in the same field. Finally, SectionĀ 6 summarizes the main findings of this work.
@@
Throughout this paper, we assume a flat CDM cosmology with = 0.3 and km/s/Mpc. Magnitudes are listed in the AB system (Oke & Gunn, 1983).
2 Data
2.1 Spectroscopic Data
@@
The sample of broad H emitters analyzed in this paper has been selected based on the spectroscopic data obtained through the JWST/NIRCam Wide Field Slitless Spectroscopy (WFSS) mode with the FRESCO and CONGRESS programs. The FRESCO survey covers 62Ā ā”ā² in each of the two GOODS fields, acquiring about 2-hour deep NIRCam/grism observations using the F444W filter. These observations provide grism spectra with a resolving power of , covering āµmĀ for most galaxies in the field-of-view (FOV). The CONGRESS covers the GOODS-N field only, with a footprint nearly identical to that of FRESCO. The observations with the F356W filter yield grism spectra with a resolving power of from 3.1 to 4.0āµm. By combining the spectroscopic data from FRESCO and CONGRESS, we are able to search for broad H emitters within a redshift range of in the GOODS-N field.
@@
All the NIRCam/WFSS data were reduced using the publicly available reduction pipeline described in Sun etĀ al. (2023). Here, we briefly summarize the main process. The WFSS data were first reduced to the stage-1 level by using the standard JWST stage-1 calibration pipeline v1.11.2. We then applied a pixel-to-pixel flat-field correction, performed the 2D sky-background subtraction, and assigned world coordinate system (WCS) information for each exposure. The WCS of the grism exposures was calibrated with the Gaia DR3 catalog (Gaia Collaboration etĀ al., 2023) by cross-matching stars identified in the simultaneously acquired NIRCam short-wavelength images with the corresponding sources in the Gaia catalog. For each source, we utilized the grism spectral tracing and dispersion models, along with the flux calibration functions presented in Sun etĀ al. (2023), to extract the 2D spectra from individual grism exposures. The extracted spectra were then coadded in a common wavelength and spatial grid. Finally, we extracted the 1D spectra using optimal extraction algorithms (Horne, 1986) and applied median filters to subtract the background and the continuum.
2.2 Multi-band Imaging Data
@@
To obtain the morphology information and estimate the physical properties of our broad H emitters, we utilize multi-band imaging data from JADES. For the GOODS-N field, 9 NIRCam images with the following filters are available: F090W, F115W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W, covering a wide range of wavelengths. However, JADES covers the GOODS-N field with a different footprint compared to those of FRESCO and CONGRESS. The total overlapping area between JADES and FRESCO/CONGRESS footprints is Ā ā”ā² in GOODS-N. Therefore, for about half of our broad H emitters, which are only covered by FRESCO and CONGRESS (see Figure 3), imaging data are available in only six filters: F090W, F115W, F182M, F210M, F356W, and F444W. Additionally, for each broad H emitter, we also checked whether deep imaging data in F435W, F606W, F775W, F814W, and F850LP are available from HST/ACS and HST/WFC3 (e.g., Koekemoer etĀ al. 2011). Overall, for our broad H emitters with , these multi-band imaging data provide excellent coverage of the rest-frame wavelengths from UV to optical (i.e., ), which allows us to constrain the physical properties of these broad H emitters through SED-fitting.
3 Broad H Emitter Selection
@@
Given the large FOV of FRESCO and CONGRESS, conducting a blind search for broad H emitters using the grism spectra across the field is extremely time-consuming. Additionally, an extended morphology can produce a fake broad-line feature in the grism data. Therefore, to identify broad H emitters more efficiently, we first construct a galaxy sample by applying compactness and photometric redshift (photo-) criteria. Specifically, we require each source to have after applying aperture correction, to ensure the selection of compact sources. To further reduce the sample size, we only include sources with photo- values reported in JADES DR2 (DāEugenio etĀ al., 2025a) between 3 and 7, which is slightly broader than the redshift range where the H emission line is expected to be observable in the grism spectra from FRESCO/CONGRESS, accounting for the uncertainty in photo-. In this way, we obtained a sample of galaxies, which is manageable for visual inspection.
@@
Next, we conducted a visual investigation of the 1D grism spectra for each galaxy to determine the presence of an H emission line and to assess whether it exhibits any broad component. For those potential broad H emitters identified by visual inspection, we fit their emission-line profiles with a multi-component Gaussian profile to determine the line widths of their broad components (see Section 3.1 for more details). To determine the origin of the broad component, we require our broad H emitters to have a broad component with a full width at half maximum (FWHM) greater than 1000 km/s. We retain only objects in which the broad component is detected with S/N 3 in line flux to ensure the broad component is real. The FWHM of the [OĀ III] doublet is also measured when available. Below, we summarize our selection criteria:
3.1 1D Line Fitting
@@
To determine the line widths of the broad component in these broad H emitters, we performed a simultaneous fit of both the narrow and broad components for H, and a narrow-component only fit for [NĀ II]6549,6585. The line ratio of the [NĀ II] doublet was fixed to 1:3.049, in accordance with recent results from DojÄinoviÄ etĀ al. (2023). Assuming that both the narrow component of H and the [NĀ II] doublet originate from the narrow line region, we constrained their line centroids and line widths to be identical. However, as the centroid of the broad H component is sometimes observed to deviate slightly from that of the narrow component (e.g., Matthee etĀ al. 2024), we allow the broad and narrow components to have different line centroids by 0.01 µmĀ (corresponding to ākm/s, depending on the redshift).
@@
In addition to fitting the line profile using a straightforward combination of narrow and broad Gaussian components, we also tested a model incorporating H absorption, although previous studies suggest that detecting such absorption features requires high S/N (e.g., Juodžbalis etĀ al. 2024; Ji etĀ al. 2025; DāEugenio etĀ al. 2025b; Rusakov etĀ al. 2025). To robustly confirm the presence of H absorption, we roughly follow the methodology outlined in Maiolino etĀ al. (2024) and perform a Bayesian Information Criterion (BIC) test (Liddle, 2007). For Gaussian noise, the BIC parameter is defined as:
@@
4 Results
4.1 19 Broad H Emitters at 3.9-5.5

### v2

@@ Abstract
We present a spectroscopic search for broad H emitters at 3.7–6.5 in the GOODS-N field, utilizing JWST/NIRCam slitless spectroscopy from FRESCO and CONGRESS, complemented by JADES imaging. We identify 19 broad H emitters with FWHM 1000 km/s at 4–5.5, including 9 new sources. The broad H luminosity function (LF) derived from our sample is consistent with those of other JWST-selected broad-line AGN reported in the literature. The black hole masses and AGN bolometric luminosities, inferred from the broad H components, indicate that most sources are accreting at 10% of the Eddington limit. We derive their host stellar masses via SED fitting and find higher ratios relative to the local – relations, consistent with previous studies. We find that 42% of the sample do not satisfy the widely-used color selection criteria for Little Red Dots (LRDs), with the majority of these sources lacking the characteristic steep rest-optical red slope, indicating that the LRD selection is highly incomplete when selecting AGN galaxies. A comparison of the average SEDs between our sample and LRDs selected in the same field reveals that the steep red slopes observed in some LRDs are likely due to line-boosting effects as previously suggested. Furthermore, we find that 68% of color-selected LRDs with H detections in the NIRCam/Grism spectra do not exhibit broad-line features. While the limited sensitivity of the grism spectra may hinder the detection of broad-line components in faint sources, our findings still highlight the enigmatic nature of the LRD population.
I Introduction
@@
Supermassive black holes (SMBHs) are thought to reside at the centers of most massive galaxies (Kormendy and Ho, 2013; Greene et al., 2020). The efficient accretion processes associated with these SMBHs contribute to the formation of active galactic nuclei (AGN), causing their host galaxies to exhibit broadened emission line profiles compared to non-active galaxies. Understanding the characteristics of AGN and accurately estimating their abundance in the early Universe () are essential for addressing key questions in extragalactic astronomy, including the reionization of the Universe and the co-evolution of AGN and their host galaxies.
@@
Prior to the launch of the JWST (Gardner et al., 2023), studying the abundance and properties of faint AGN at high redshifts was highly challenging due to observational limitations. For AGN with ultraviolet (UV) magnitudes fainter than , determining their space density and physical properties was particularly difficult, with significant discrepancies reported (e.g., Parsa et al., 2018; Giallongo et al., 2019; Morishita et al., 2020; Shen et al., 2020; Finkelstein and Bagley, 2022). The advent of JWST, with its unparalleled infrared sensitivity, has opened new avenues for investigating these faint, distant objects. Recent JWST observations have demonstrated its remarkable capability to identify UV-faint AGN in the early Universe through various methodologies, including the detection of broad Balmer lines (Kocevski et al., 2023; Harikane et al., 2023; Oesch et al., 2023; Larson et al., 2023; Übler et al., 2023; Barro et al., 2024; Matthee et al., 2024; Maiolino et al., 2024; Greene et al., 2024; Lin et al., 2025d).
@@
Among the UV-faint AGN detected by JWST, an intriguing subset of the AGN population consists of sources with notably compact sizes and red colors. Matthee et al. (2024) assembled one of the earliest samples of these sources, identified through broad H emission lines, and named them as “little red dots” (LRDs). These LRDs appear as point sources and exhibit distinctive spectral energy distributions (SEDs), characterized by blue rest-frame UV slopes and extremely steep red rest-frame optical slopes (e.g., Greene et al., 2024; Labbé et al., 2025; Ji et al., 2025; Naidu et al., 2025; de Graaff et al., 2025). Such a “V-shaped” SED is found to be rare for AGN at 2 (e.g., Noboriguchi et al. 2019). To investigate the origin of the UV and optical emission in LRDs, several recent studies have assembled samples of these objects by identifying compact sources and targeting those with “V-shaped” SEDs, using NIRCam color cuts as a selection criterion (Akins et al., 2023; Barro et al., 2024; Greene et al., 2024; Pérez-González et al., 2024; Kokorev et al., 2024; Williams et al., 2024; Rinaldi et al., 2024; Labbé et al., 2025).
@@
Although numerous photometrically selected LRD candidates are assembled, their true nature remains uncertain. It is still unclear whether these LRDs are primarily AGN, in which the red optical continua arise from dust emission in front of UV-bright AGN (e.g., Akins et al., 2023; Maiolino et al., 2024) or non-stellar Balmer breaks (Lin et al., 2025b), or whether they are dusty star-forming galaxies, where strong emission lines such as H or [O III] with high equivalent widths can boost the broad-band fluxes and make the rest-frame optical colors appear red (e.g., Barro et al., 2024; Pérez-González et al., 2024; Hviding et al., 2025). An additional possibility is that some LRDs may be “black hole stars” embedded in relatively bright host galaxies, in which extremely dense gas forms a dust-free atmosphere around a SMBH, producing strong Balmer break and absorption features (Naidu et al., 2025). Greene et al. (2024) obtained follow-up NIRSpec spectroscopy on 15 LRDs photometrically selected in Labbé et al. (2025) and found that 60% exhibited definitive evidence of broad H lines, confirming their AGN nature for a significant subset. In contrast, Pérez-González et al. (2024) reported that only 17% of their photometrically selected LRDs with available spectroscopy displayed broad line features. However, this fraction may be underestimated, as many existing broad lines could not be confirmed due to insufficient S/N of the grism spectra (Williams et al., 2024). In Rinaldi et al. (2024), analysis of the emission line properties of the LRD sample using NIRSpec spectra and classical line ratio diagnostics was unable to conclusively determine their nature, suggesting a likely mixed origin.
@@
In this work, we utilized grism spectroscopic data from the First Reionization Epoch Spectroscopically Complete Observations (FRESCO-PID: 1895; PI: P. Oesch; Oesch et al. 2023) and the Complete NIRCam Grism Redshift Survey (CONGRESS-PID: 3577; PI: E. Egami; Sun et al., in prep, Lin et al. 2025a), along with imaging data from the JWST Advanced Deep Extragalactic Survey (JADES-PID: 1181; PI: D. Eisenstein; Eisenstein et al. 2023; Rieke, Marcia et al. 2023), to conduct a comprehensive search for broad H emitters at in the GOODS-North field (Giavalisco et al. 2004, hereafter GOODS-N). After identifying this sample of broad H emitters, we aim to characterize their physical properties and compare them with LRDs recently identified in other studies. The environments of these AGNs and their impact on BH growth are discussed in a companion paper (Lin et al., 2025c).
@@
This paper is organized as follows: In Section II, we provide a brief overview of the spectroscopic and imaging data utilized in this study. Section III details the selection process for broad H emitters and the methodology for 1D line fitting. In Section IV, we present and discuss the results of the 1D line fitting for our broad H emitters. Section V explores the physical properties of the broad H emitter sample and compares our sample with the LRD sample identified in the same field. Finally, Section VI summarizes the main findings of this work.
@@
Throughout this paper, we assume a flat CDM cosmology with = 0.3 and km/s/Mpc. Magnitudes are listed in the AB system (Oke and Gunn, 1983).
II Data
II.1 Spectroscopic Data
@@
The sample of broad H emitters analyzed in this paper has been selected based on the spectroscopic data obtained through the JWST/NIRCam Wide Field Slitless Spectroscopy (WFSS) mode with the FRESCO and CONGRESS programs. The FRESCO survey covers 62 □′ in each of the two GOODS fields, acquiring about 2-hour deep NIRCam/grism observations using the F444W filter. These observations provide grism spectra with a resolving power of , covering µm for most galaxies in the field-of-view (FOV). The CONGRESS covers the GOODS-N field only, with a footprint nearly identical to that of FRESCO. The observations with the F356W filter yield grism spectra with a resolving power of from 3.1 to 4.0 µm. Since the GOODS-S field has grism spectroscopic coverage only in F444W, and Matthee et al. (2024) have already performed a systematic search for broad H emitters using these data, this study focuses on the GOODS-N field only. By combining the spectroscopic data from FRESCO and CONGRESS, we are able to search for broad H emitters within a redshift range of in the GOODS-N field.
@@
All the NIRCam/WFSS data were reduced using the publicly available reduction pipeline described in Sun et al. (2023). Here, we briefly summarize the main process. The WFSS data were first reduced to the stage-1 level by using the standard JWST stage-1 calibration pipeline v1.11.2. We then applied a pixel-to-pixel flat-field correction, performed the 2D sky-background subtraction, and assigned world coordinate system (WCS) information for each exposure. The WCS of the grism exposures was calibrated with the Gaia DR3 catalog (Gaia Collaboration et al., 2023) by cross-matching stars identified in the simultaneously acquired NIRCam short-wavelength images with the corresponding sources in the Gaia catalog. For each source, we utilized the grism spectral tracing and dispersion models, along with the flux calibration functions presented in Sun et al. (2023), to extract the 2D spectra from individual grism exposures. The extracted spectra were then coadded in a common wavelength and spatial grid. Finally, we extracted the 1D spectra using optimal extraction algorithms (Horne, 1986) and applied median filters to subtract the background and the continuum.
II.2 Multi-band Imaging Data
@@
To obtain the morphology information and estimate the physical properties of our broad H emitters, we utilize multi-band imaging data from JADES. For the GOODS-N field, 9 NIRCam images with the following filters are available: F090W, F115W, F150W, F200W, F277W, F335M, F356W, F410M, and F444W, covering a wide range of wavelengths. However, JADES covers the GOODS-N field with a different footprint compared to those of FRESCO and CONGRESS. The total overlapping area between JADES and FRESCO/CONGRESS footprints is □′ in GOODS-N. Therefore, for about half of our broad H emitters, which are only covered by FRESCO and CONGRESS (see Figure 3), imaging data are available in only six filters: F090W, F115W, F182M, F210M, F356W, and F444W. Additionally, for each broad H emitter, we also checked whether deep imaging data in F435W, F606W, F775W, F814W, and F850LP are available from HST/ACS and HST/WFC3 (e.g., Koekemoer et al. 2011). Overall, for our broad H emitters with , these multi-band imaging data provide excellent coverage of the rest-frame wavelengths from UV to optical (i.e., ), which allows us to constrain the physical properties of these broad H emitters through SED-fitting.
III Broad H Emitter Selection
@@
Given the large FOV of FRESCO and CONGRESS, conducting a blind search for broad H emitters using the grism spectra across the field is extremely time-consuming. Additionally, an extended morphology can produce a fake broad-line feature in the grism data. Therefore, to identify broad H emitters more efficiently, we first construct a galaxy sample by applying compactness and photometric redshift (photo-) criteria. Specifically, we require each source to have after applying aperture correction, to ensure the selection of compact sources. To further reduce the sample size, we only include sources with photo- values reported in JADES DR2 (D’Eugenio et al., 2025a) between 3 and 7, which is slightly broader than the redshift range where the H emission line is expected to be observable in the grism spectra from FRESCO/CONGRESS, accounting for the uncertainty in photo-. In this way, we obtained a sample of galaxies, which is manageable for visual inspection.
@@
Next, we conducted a visual investigation of the 1D grism spectra for each galaxy to remove contaminants that do not show a clear H emission line, either because no line is detected or because the observed line is inconsistent with H. For those potential broad H emitters identified by visual inspection, we fit their emission-line profiles with a multi-component Gaussian profile to determine the line widths of their broad components (see Section III.1 for more details). To select sources with broad-line feature, we require our broad H emitters to have a broad component with a full width at half maximum (FWHM) greater than 1000 km/s. We retain only objects in which the broad component is detected with S/N 3 in line flux to ensure the broad component is real. The FWHM of the [O III] doublet is also measured when available. Below, we summarize our selection criteria:
III.1 1D Line Fitting
@@
To determine the line widths of the broad component in these broad H emitters, we performed a simultaneous fit of both the narrow and broad components for H, and a narrow-component only fit for [N II]6549,6585. The line ratio of the [N II] doublet was fixed to 1:3.049, in accordance with recent results from Dojčinović et al. (2023). Assuming that both the narrow component of H and the [N II] doublet originate from the narrow line region, we constrained their line centroids and line widths to be identical. However, as the centroid of the broad H component is sometimes observed to deviate slightly from that of the narrow component (e.g., Matthee et al. 2024), we allow the broad and narrow components to have different line centroids by 0.01 µm (corresponding to km/s, depending on the redshift).
@@
In addition to fitting the line profile using a straightforward combination of narrow and broad Gaussian components, we also tested a model incorporating H absorption, although previous studies suggest that detecting such absorption features requires high S/N (e.g., Juodžbalis et al. 2024; Ji et al. 2025; D’Eugenio et al. 2025b; Rusakov et al. 2025). To robustly confirm the presence of H absorption, we roughly follow the methodology outlined in Maiolino et al. (2024) and perform a Bayesian Information Criterion (BIC) test (Liddle, 2007). For Gaussian noise, the BIC parameter is defined as:
@@
IV Results
IV.1 19 Broad H Emitters at 3.9-5.5